OpenStudy (anonymous):

Differential equations? Calculus? solve the differential equation. f"(s) = sin x, f'(0)=1, f(0)=6

OpenStudy (unklerhaukus):

do you mean f"(x)?

OpenStudy (anonymous):

yes sorry

OpenStudy (mathmale):

Good morning, VBG! Something seems not quite right here. If f ' (x)=sin x and f ' (0), that appears to be two different definitions. Am I correct in supposing you meant f '' (x) = sin x?

OpenStudy (mathmale):

Assuming that It's f ' '(x) that equals sin x, then f ' (x) = -cos x +C and f '(0) = 1 = -cos 0 + C. Make sense?

OpenStudy (unklerhaukus):

(laplace method is more fun)

OpenStudy (mathmale):

Surely is. Vilbarguen is currently offline.

OpenStudy (unklerhaukus):

i hadn't thought of applying those initial conditions mid-way (so-to-speak) through solving, i usually wait till the end, but i see that will work fine and even make the method simpler

OpenStudy (anonymous):

Hey guys. Sorry I had to leave. yes it is f '' (x), i just used " so I guess it looked like ' .

OpenStudy (anonymous):

mathmale so f ' (x) = -cos x +C because it is the antiderivative yes?

OpenStudy (mathmale):

Yes, i was confused there. But that's not your fault. Integrate f ' ' (x) and add the constant of integration, C. Find C, based upon the initial conditions.

OpenStudy (mathmale):

Your f ' (x" is fine. Can you now find C? I know you must be working on that right now.

OpenStudy (anonymous):

so if f ' (0)= 1, I use - cos (0) + C = -1 + C = -1 + 2 f'(0) = 1

OpenStudy (anonymous):

am I right?

OpenStudy (mathmale):

You are saying that C=2, right? If so, then take your f '(x)=-cos x +2 and integrate every term with respect to x. Add another constant of integration, D.

OpenStudy (anonymous):

hold on I don't understand. this is to solve f(0)=6 ?

OpenStudy (mathmale):

VBG: You are to integrate your expression for f ' (x), adding a new constant, D, at the end. Once you've done that, you take the last initial condition and insert those values into your f(x). Here's what that would look like: f(0)=6=-sin 0 +2(0) +D. Find D. then write

OpenStudy (mathmale):

f(x) = ???? incorporating C=2 and D = ? and then you're done.

OpenStudy (anonymous):

f(0)= 6= -sin (0) + 2 (0) + D = 0 + 0 + D=6 f(0) = 6

OpenStudy (anonymous):

thats it

OpenStudy (mathmale):

So, f(x) = ???

OpenStudy (mathmale):

VBG: Please summarize your result: f(x) = ?? If you're tied up with something else, or prefer not to do this, please just say so.

OpenStudy (anonymous):

so f(x) = -sin (x) + 2(x) + D ?

OpenStudy (anonymous):

sorry about the delay

OpenStudy (anonymous):

D= 6 and C= 2 for f'(x)

OpenStudy (mathmale):

Yes, and y ou've found that D=6 so you end up with f(x)=-sin x +2x +6, right? You could check this by differentiating, if you want to. Any other concerns related to this homework problem?

OpenStudy (anonymous):

ah ok so I put 6 instead of D. got it.

OpenStudy (anonymous):

No mathmale thats it. Thank you so much for your help

OpenStudy (mathmale):

You're very welcome. De nada. Hope to "meet" with you on OpenStudy again soon. Great day to you. Bye.

OpenStudy (anonymous):

Certainly. Thanks again!